HIGH-ACCURACY DIABATIC TREATMENT OF NO_{3} ENERGY LEVELS

\underline{JOHN F. STANTON}, Department of Chemistry & Biochemistry, The University of Texas at Austin, 1 University Station A5300 Austin, TX 78712.

Recent developments in vibronic coupling theory and quantum chemistry now allow the parametrization of vibronic Hamiltonians that are capable of achieving semi-quantitative agreement with experimental energy levels for strongly coupled systems. First applied to the HCO_{2} and BNB radicals, this sophisticated parametrization has been used for the nitrate radical. Results of calculations bearing on both all electronic states of NO_{3} below 2 eV will be presented.